# Regret Optimal Control for Uncertain Stochastic Systems

### v2

@@ V Conclusion
We have presented a novel method for convex synthesis of robust control policies with provable regret and safety guarantees in face of the uncertain stochastic dynamics. As the clairvoyant optimal policy we compete against is unknown in this setting, we have proposed sampling the space of parameters that characterize the system dynamics. Leveraging results from the theory of scenario optimization, we have shown that the policy that minimizes regret robustly over these randomly drawn uncertainty instances retains strong probabilistic out-of-samples guarantees. Finally, we have presented numerical experiments to corroborate our theoretical results, and to highlight the potential of regret minimization in adapting to heterogeneous dynamics and disturbance sequences. Interesting directions for future research encompass studying infinite-horizon control problems, addressing computational complexity challenges for real-time implementation, and extending the theory of this emerging competitive framework to systems with nonlinear dynamics.

### v3

@@ 5 Conclusion
We have presented a novel method for convex synthesis of robust control policies with provable regret and safety guarantees in face of the uncertain stochastic dynamics. As the clairvoyant optimal policy we compete against is unknown in this setting, we have proposed sampling the space of parameters that characterize the system dynamics. Leveraging results from the theory of scenario optimization, we have shown that the policy that minimizes regret robustly over these randomly drawn uncertainty instances retains strong probabilistic out-of-samples guarantees. Finally, we have presented numerical experiments to corroborate our theoretical results, and to highlight the potential of regret minimization in adapting to heterogeneous dynamics and disturbance sequences. Interesting directions for future research encompass studying infinite-horizon control problems, devising novel solutions that do not rely on sampling in order to robustify against dynamic and non-parametric uncertainties, addressing computational complexity challenges for real-time implementation, and extending the theory of this emerging competitive framework to systems with nonlinear dynamics.
Appendix A Proof of Proposition 1
For any $\bm{\theta}^{k}\in\mathcal{D}$, by combining (2) with (7), we first rewrite the per-instance regret (3) as
$\left\lVert\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Phi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Phi}_{u}\end{bmatrix}\bm{w}\right\rVert^{2}_{2}-\left\lVert\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}\bm{w}\right\rVert^{2}_{2}\,.$ (18)
By gathering common terms, (18) can be equivalently expressed as the quadratic form $\bm{w}^{\top}\bm{\Delta}(\bm{\Phi}_{u},\bm{\Psi}_{u}(\bm{\theta}^{k}),\bm{\theta}^{k})\bm{w}$, where $\bm{\Delta}(\bm{\Phi}_{u},\bm{\Psi}_{u}(\bm{\theta}^{k}),\bm{\theta}^{k}):=\bm{\Delta}^{k}$ is defined as
${\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Phi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Phi}_{u}\end{bmatrix}}^{\top}\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Phi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Phi}_{u}\end{bmatrix}-{\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}}^{\top}\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}.$
From well-known properties of induced matrix norms and from classical results on semidefinite programming for eigenvalue minimization (see, e.g., Section 2.2 in [45]), we have:
$\max_{\|\bm{w}\|_{2}\leq 1}~{}\bm{w}^{\top}\bm{\Delta}^{k}\bm{w}=\lambda_{\max}(\bm{\Delta}^{k})=$ $~{}\min_{\lambda^{k}}~{}\lambda^{k}$
$\operatorname*{subject~{}to}~{}\lambda^{k}\bm{I}\succeq\bm{\Delta}^{k}\,.$
In other words, for a fixed $\bm{\theta}^{k}\in\mathcal{D}$, the regret bound (10d) is satisfied if and only if the robust performance level $\gamma$ is not smaller than the minimum $\lambda^{k}$ such that $\lambda^{k}\bm{I}-\bm{\Delta}^{k}\succeq 0$. Equivalently, but without introducing an unnecessary optimization variable for each sampled scenario, if and only $\gamma\bm{I}-\bm{\Delta}^{k}\succeq 0$. From this last expression, the set of linear matrix inequality constraints (11c) follows by exploiting the Schur complement to remove the quadratic dependence of $\bm{\Delta}^{k}$ on $\bm{\Phi}_{u}$.
We then turn our attention to enforcing robust satisfaction of the safety constraints (10c). First, we note that each row $\max_{\bm{w}\in\mathcal{W}(\bm{\theta}^{k})}~{}(\bm{H}_{x}^{k}\bm{\Phi}_{x}(\bm{\theta}^{k})+\bm{H}_{u}^{k}\bm{\Phi}_{u})_{i}\bm{w}$ of (10c) constitutes a second-order cone program in $\bm{d}$, where $\bm{d}$ is such that $\bm{w}=\bm{H}_{w}(\bm{\theta}^{k})\bm{d}$ as per (5). Leveraging well-known properties of dual norms, we then observe that $\max_{\left\lVert\bm{d}\right\rVert_{2}\leq 1}~{}(\bm{H}_{x}^{k}\bm{\Phi}_{x}(\bm{\theta}^{k})+\bm{H}_{u}^{k}\bm{\Phi}_{u})_{i}\bm{H}_{w}^{k}\bm{d}$ is equal to
$\left\lVert{\bm{H}_{w}^{k}}^{\top}(\bm{H}_{x}^{k}\bm{\Phi}_{x}(\bm{\theta}^{k})+\bm{H}_{u}^{k}\bm{\Phi}_{u})_{i}^{\top}\right\rVert_{2}\,.$
Lastly, iterating over all rows of (10c), the set of second-order cone constraints (11b) follows.
Appendix B Proof of Theorem 1
In light of Proposition 1, we have that the real-valued functions $f_{1}(\bm{\Phi}_{u},\gamma,\bm{\theta})=\max_{\left\lVert\bm{w}\right\rVert_{2}\leq 1}~{}\mathtt{R}(\bm{\Phi}_{u},\bm{\Psi}_{u}(\bm{\theta}),\bm{w},\bm{\theta})-\gamma$ and $f_{2}(\bm{\Phi}_{u},\bm{\theta})=\max_{\bm{w}\in\mathcal{W}(\bm{\theta})}~{}\bm{H}_{x}(\bm{\theta})\bm{\Phi}_{x}(\bm{\theta})\bm{w}+\bm{H}_{u}(\bm{\theta})\bm{\Phi}_{u}\bm{w}-\bm{h}(\bm{\theta})$, with $\bm{\Phi}_{x}(\bm{\theta})$ a linear map of $\bm{\Phi}_{u}$ as per (7), are convex in the design parameters $\bm{\Phi}_{u}$ and $\gamma$. Since the point-wise maximum of convex functions is convex, we then conclude that $f(\bm{\Phi}_{u},\gamma,\bm{\theta})=\max~{}(f_{1}(\bm{\Phi}_{u},\gamma,\bm{\theta}),f_{2}(\bm{\Phi}_{u},\bm{\theta}))$ is also convex in $\bm{\Phi}_{u}$ and $\gamma$. Hence, the robustly safe regret minimization problem (9) effectively calls for the minimization of a linear objective subject to a possibly infinite number of convex constraints of the form $f(\bm{\Phi}_{u},\gamma,\bm{\theta})\leq 0$. Based on these observations, our probabilistic guarantees then follow by combining the results of Theorem 1 in [19] and Theorem 2.4 in [20].